I am having trouble solving this eqation... x/6 + 3x/4 +1 =2x/3 + 5/6
Answer by stanbon(75887)

x/6 + 3x/4 + 1= 2x/3 + 5/6
------------------
Multiply thru by 12 to get:
---
2x + 3*3x + 12 = 4*2x + 2*5
----
2x+9x +12 = 8x+10
11x + 12 = 8x + 10
3x = -2
x = -2/3
===================
---------------------------
Cheers,
Stan H.